Circular No. 5892

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET SHOEMAKER-LEVY 9 (1993e) J. V. Scotti and T. Metcalfe, Lunar and Planetary Laboratory, have

provided accurate absolute astrometry for nine of the nuclei of this comet from CCD images obtained with the 0.91-m Spacewatch telescope at Kitt Peak generally on nine nights during Mar. 30-July 19. The observations and the orbits determined from them by the undersigned are being published in the Nov. 29 batch of Minor Planet Circulars. The nuclei involved are those numbered 1, 5, 6, 7, 11, 12, 14, 15 and 17 by Jewitt et al. (1993, Bull. Am. Astron. Soc. 25, 1042), and the orbit for the brightest nucleus (nucleus 7) is as follows:

Epoch = 1994 May 8.0 TT

T = 1994 Apr. 1.3226 TT  ω = 355.^{o}1307
e = 0.206613  Ω = 220.9658 }2000.0
q = 5.382476 AU  i = 5.7864

a = 6.784178 AU n^{o} = 0.0557774 P = 17.670 years It must be emphasized that these computations were made for the

various nuclei independently. All the orbits agree in giving a minimum distance of 0.0008 AU from the center of Jupiter on 1992 July 8.0 UT (cf. IAUC 5800), but the computed sky-plane positions of the nuclei at that time understandably differ (by up to 12′′ in α and 16′′ in δ, with the poorly observed nucleus 6 showing a shift of up to 70′′ in δ)—although the orbits placed nuclei 5, 11 and 17 then within a region only 1′′.2×1′′.7. Nevertheless, the close agreement of the ephemerides extrapolated back to the beginning of 1992, and even to the previous perijove 0.06 AU from Jupiter in July 1989 (with the indication that there were several earlier revolutions around Jupiter), clearly demonstrates that the basic break-up occurred very close to the 1992 perijove and involved very small velocity differences.

The new computations indicate that the 1992 prediscovery ephemeris on IAUC 5801 requires adjustment as follows: Feb. 8, Δα = −1^{m}.13, Δδ = −12′.6; Mar. 9, −0^{m}.81, −9′.7; Apr. 8, −0^{m}.51, −6′.5. G. Tancredi, University of Montevideo; and M. Lindgren and C.-I. Lagerkvist, Uppsala University, report that 90-min Kodak IIIa-J exposures of the region in Mar. 1992 with the 1.0-m Schmidt telescope at the European Southern Observatory and tracked at the motion of Jupiter showed no trace of the comet. On 1992 Mar. 6 the limiting magnitude for a stationary object was B = 22.1 (or B = 21.3 for an object having the comet’s motion).

1993 November 22 Brian G. Marsden